# VideoHallu: Evaluating and Mitigating Multi-modal Hallucinations for Synthetic Videos

### v1

@@ Abstract
Synthetic video generation using foundation models has gained significant attention due to its realism and broad applications. However, while these models excel at generating visually coherent and high-quality video frames, they often overlook commonsense reasoning and physical law violations, leading to abnormal content.
Existing score-based evaluations like VideoScore [1] mainly focus on general video quality and do not take these abnormalities into account, and offer no explanations of the evaluation results. A more promising evaluation approach is to leverage multi-modal large language models (MLLMs) as interpretable video evaluators, following the approach of FactScore [2]. However, how well MLLMs can detect these abnormalities in synthetic videos is underexplored. Motivated by a more interpretable video generation evaluation, we introduce VideoHallu, a benchmark built from synthetic videos produced by popular models like Veo2 [3], Sora [4], and Kling [5], paired with expert-crafted question-answering pair examples easily solvable with human-level perception and reasoning across multiple categories. We evaluate several State-of-the-Art (SoTA) MLLMs with our benchmark, including GPT-4o [6], Gemini-2.5-Pro [7], Qwen-2.5-VL [8], and forefront models like Video-R1 [9] and VideoChat-R1 [10]. Despite the strong performance of R1 MLLMs on real-world video benchmarks like MVBench [11] and MovieChat [12], these models still struggle and hallucinate on basic commonsense and physics reasoning tasks in synthetic videos, highlighting synthetic video hallucination as an underexplored challenge. Moreover, we post-train current SoTA MLLMs with Group Relative Policy Optimization (GRPO) [13] using both real-world and synthetic commonsense/physics datasets. Our results show improved overall accuracy compared to the base model, achieving the highest performance among all models, highlighting the importance of integrating high-quality counterexamples to enhance commonsense and physics reasoning in MLLMs’ language priors. Our data is available at https://github.com/zli12321/VideoHallu.git.
@@ 1 Introduction
Figure 1: Hallucinations in Synthetic Video Understanding. Synthetic videos often exhibit counterfactual or common-sense-violating behaviors due to current video generation model limitations. While prior evaluation works mainly score video quality and consistency, our work, VideoHallu, probes hallucinations caused by misapplied common sense in video-dependent questions. We design both common sense QA (answerable without video) and video-dependent QA (requiring video context). Evaluations show that SoTA MLLMs often hallucinate on video-dependent QA, relying on language priors instead of video inputs.
Foundational video generation models [14] have attracted significant attention due to their recent success in AI-Generated Content (AIGC). Their realistic and high-fidelity outputs open broad applications across multimedia [15, 16], entertainment and content creativity [17], robotics and embodied AI [18]. Despite the promise of current video generation models, many critical flaws emerge during deployment. A key limitation lies in their failure to faithfully generate reality-grounded videos beyond following basic input prompt instructions and enhancing general video quality such as resolution [1]. Many state-of-the-art (SoTA) video generation models like Gemini-2.5-Pro [7], GPT-4o [19] and Qwen2.5-VL [8] (Figure. 1) exhibit hallucination issues such as inconsistent motion dynamics [20], and violations of real-world common-sense [21] and basic physics laws [22, 23]. These shortcomings suggest that video generation models mainly optimize input-output alignment, reflected in the loss designs of CogVideo [24] and LaVie [25], without truly learning or applying real-world commonsense and physical laws. Instead, they tend to superficially imitate training data appearance and dynamics, rather than fundamentally understand the underlying principles of real-world behavior.
Detecting problematic generations make evaluating video generation more crucial, especially at scale [23, 26]. Socre-based evaluation methods such as VideoScore [1] or DEVIL [27] produce a set of video quality scores without explanations, and they suffer from out-of-distribution data (newer released video generation models) and do not align with human visual judgments, as Figure. 1 suggests. One promising approach to facilitate a more interpretable video generation evaluation is by leveraging MLLMs for synthetic video understanding and evaluation. MLLMs offer insights into video understanding through their grasp of driving principles and commonsense knowledge, and can detect abnormalities by breaking down input prompts into subsets of questions and individually querying the video like FactScore [2] style, facilitating a more interpretable evaluation of video generation. However, current MLLMs are highly prone to hallucinations when evaluating synthetic videos. As shown in Figure. 1, even SoTA models hallucinate when presented with synthetic videos that are visually flawed and counter-intuitive, despite their strong performance on real-world videos [11, 28]. We attribute this to the scarcity of high-quality and high-fidelity synthetic annotated videos in previous MLLMs’ training data, as most are trained with real-world video data source [29, 8, 7, 30]. Due to the lack of synthetic videos in the training data, MLLMs can only rely on their LLM backbones to reason, treating commonsense and physical knowledge as default ground truth assumptions without considering the actual visual input [31], which introduces biases stemming from the LLM’s strong prior knowledge. While these biases may not surface on real-world videos, they cause severe hallucinations on synthetic videos, which often contain counterintuitive phenomenons that violate commonsense and physical knowledge (Figure. 1). Furthermore, although chain-of-thought reasoning often mitigates hallucinations in real-world understanding [32], the reasoning itself inherits deep-rooted hallucinations within MLLMs’ language priors due to the scarcity of annotated negative examples in synthetic video understanding tasks. Current MLLMs are unsuitable for reliably detecting commonsense violations in synthetic videos.
@@
We propose VideoHallu, a benchmark consisting of expert-crafted, diverse question-answering pairs spanning context alignment, consistency, commonsense, and physical law reasoning of synthetically generated videos to bridge the gap between video generation and synthetic video evaluation. These reasoning-heavy tasks require an in-depth understanding of visual cues and language-based knowledge to answer our questions. We evaluate several SoTA video understanding models and provide a detailed analysis of their current failures on synthetic video understanding. Additionally, we use curriculum learning [33] and Group Relative Policy Optimization (GRPO) [13] to post-train an MLLM using datasets combining real-world (VideoLLaVA [34], PhysBench [22]) and synthetic videos (VideoHallu) to validate the effectiveness of improving reasoning abilities on synthetic video tasks through post-training. Our goal is to improve commonsense and physical law reasoning into synthetic video understanding. We hope VideoHallu and our RFT (Reinforcement Fine Tuning) framework offer insights into mitigating hallucinations and advancing video understanding and evaluation for synthetic videos. Our contributions include:
@@
  • We propose a novel benchmark VideoHallu, to test synthetic video understanding and evaluation tasks, featuring expert-annotated, diverse question-answer pairs covering alignment, spatial-temporal consistency, commonsense, and physical reasoning tasks.
  • We conduct evaluation experiments on several state-of-the-art video understanding models using VideoHallu. Results show that the top-performing models achieve only a modest overall accuracy of 51.4%, with a highly uneven performance across perceptual-heavy and reasoning-heavy questions—exhibiting a gap of over 20% accuracy between the two.
@@
  • We use curriculum learning and GRPO to fine-tune QWen2.5-VL-7B using a small subset in combination of real-world and synthetic video understandings–Video-LLaVA-80K [34], Physbench [22], and VideoHallu. Our results show that even a limited amount of RFT synthetic data (800) yields a notable 3% overall accuracy improvement, with significant gains in commonsense and physics reasoning question accuracy, highlighting its strong potential for future exploration.
2 Related Work
2.1 Hallucination in Vision-language Models
Hallucinations in MLLMs—i.e., the tendency to generate outputs that are inconsistent with the target texts, images, or videos [35, 36]—remain a long-standing challenge. Previous works have focused on developing benchmarks or pipelines to trigger hallucinations [31, 37, 38], as well as mitigating them through instruction tuning [39] or by enhancing the model’s visual and textual representations [40]. Since hallucination arises from conflicts between the language priors of MLLMs and the actual visual inputs [38], it can be even more severe in video understanding than in static image understanding—due to the complex entanglement of spatial-temporal information across the timeline and the contextual cues associated with entities within frames. A line of prior work, such as VideoHallucer[41] and EventHallusion[42], has aimed to establish benchmarks for evaluating model hallucination on both entities and events within videos, while also proposing methods to enhance the video understanding capabilities of MLLMs. HAVEN [43] further investigates the causes of hallucinations in video understanding models and introduces a video-thinking framework that incorporates reasoning and reflective thinking to mitigate hallucinations. However, most prior works on hallucination—particularly in the video domain—rely on real-world factual data, rather than synthetic data generated by generative models. Hallucination in generative video understanding models remains an open and largely unexplored research area.
2.2 Video Understanding Models
Video understanding is a fundamental task in computer vision [44, 45]. Traditional video understanding [46, 47] focuses on understanding the events happening over the timeline throughout the video. The emergence of foundation models in video understanding tasks [48, 49] has enabled the development of general-purpose models capable of understanding videos across diverse domains, by leveraging large-scale training data and knowledge encoding through increased model capacity. Prior works mainly focus on using the instruction tuning framework [50] to unify the representation of video and language prompt into the same representation space [34, 30] or scale up the visual-language models into the video domain [8, 19, 7]. Given the diverse deployment scenarios for video understanding models, recent works have focused on designing or distilling lightweight expert models for tasks such as video question answering [51, 52, 53]. These models aim to enable applications on edge devices and have achieved competitive performance compared to large-scale models while using only a fraction of the parameters. Meanwhile, rule-based reinforcement learning has also been applied during the fine-tuning process [9, 10] to encourage models to exhibit self-improvement and chain-of-thought reasoning abilities in video understanding. Despite the success of existing large models in real-world video understanding, their limitations—such as poor performance in probing commonsense and physical law violations in videos, as well as hallucinations in question answering—continue to hinder their broader adoption across diverse application scenarios. [54].
2.3 Evaluation of Video Generation Models
Video generation models [3, 5, 55, 25, 56, 24] have a wide range of applications, including content creation, robotic environment simulation, and training [49]. However, the limited performance of video generation models—characterized by cross-modal misalignment [21] and frequent violations of commonsense [22]—remains a key bottleneck to their broader adoption and application. Many efforts have been devoted to designing comprehensive benchmarks for evaluating the quality of generated videos [57, 58], as well as developing model judges [59] to assess the fidelity and coherence of generated content. Early explorations have primarily focused on incorporating human perception and preference annotations into the evaluation of video generation models [21, 60], or on developing versatile evaluators that assess multiple dimensions of video quality, including spatial-temporal consistency, cross-modal alignment, and visual fidelity [1]. As video generation models continue to advance, recent efforts have shifted toward context-level evaluation of dynamic properties and actions, with a growing emphasis on assessing the representation of commonsense knowledge implicitly embedded in generation prompts [61], particularly those related to physical laws [62, 63]. Notable investigations in this direction leverage human preference data to fine-tune evaluation models for detecting violations of physical laws arising from entity motion [23], or broader commonsense violations in layout generation and state transitions within large-scale scenes or dynamic entity interactions, often through the design of expert-informed metrics [64]. Despite substantial progress in this field, many open questions remain, as prior work has largely focused on quantitative evaluation of video quality rather than offering interpretable assessments or insights into entity-wise motion understanding and reasoning over evaluation outcomes—gaps that this work aims to address.
@@ 3 VideoHallu: Benchmark for Synthetic Video Understanding and Evaluation
3.1 Preliminary
Synthetic Videos.
This work focuses on video understanding tasks using synthetic videos—i.e., videos generated by foundation models from textual prompts. Ideally, such videos should not only align with the prompts but also exhibit real-world plausibility, including: (1) smooth temporal changes in entities, (2) object geometry and dynamics consistent with common sense, and (3) entity motion that adheres to physical laws unless stated otherwise. We manually curate prompts to evaluate how well video generation models align with our predefined categories. For each prompt, we generate videos using seven SoTA models: Sora [4], Veo2 [3], Kling [5], Runway Gen2 [56], PixVerse [65], LaVie [25], and CogVideo [24].
An example prompt for our synthetic video generation is:
Generate a scene of a feather and a heavy rock are released at the same height and begin to fall to the ground on Earth.
We keep prompts simple and specific to avoid confusing the generation models.
Synthetic Video Understanding and Evaluation.
Despite growing interest in video generation, synthetic videos generated by current foundation models remain far from flawless. Visual abnormalities and counterintuitive phenomena are common, underscoring the need for models that can both evaluate video quality and detect such anomalies. MLLMs are emerging as prominent candidates for serving as comprehensive video judges. However, beyond basic quality assessment or scene description, synthetic video understanding requires models to: (1) detect, localize, and characterize anomalies; (2) apply commonsense and physical reasoning to assess plausibility; and (3) resist hallucinations when visual input conflicts with their language priors. This threefold challenge sets a higher standard for MLLMs, requiring them to approach human-level video perception to perform effectively.
3.2 Case study: Can MLLMs understand complex concepts as humans do in synthetic videos, and are current video evaluation models truly reliable?
A natural approach to using MLLMs as video judges, which can be done using either score-based models, VideoScore [1] and WorldScore [64] that assign scores to aspects like visual quality, alignment with prompts, and factual consistency, or prompt video understanding MLLMs with target-specific questions and performing answer matching or directly give a quality rating score [26]. To assess whether both kinds of models can handle synthetic video understanding tasks with awareness of anomalies, we present a case study as part of our assessment:
We first generate a synthetic video by prompting two video generation models, Veo2 [3] and Runway Gen2 [56], as follows
A feather and a heavy rock are released into the air and start to fall.
As shown in Figure 1, the Veo2-generated video exhibits a commonsense violation, while the Runway Gen2 video is misaligned with the prompt and lacks physical consistencies. This case is designed to evaluate whether video judge models can resolve conflicts between their language priors and actual visual observations.
We first apply VideoScore [1] to assess both videos across metrics like visual quality, text-video alignment, and consistency. Both videos receive similar, modest scores across all metrics. However, the misalignment in the Runway Gen2 video is clearly noticeable, suggesting a gap in VideoScore’s ability to detect such issues, which is consistent to conclusion [26] that MLLMs are still far from being as qualified video judges than humans do. Meanwhile, Veo2 video contains a commonsense violation not covered by VideoScore’s evaluation scope, which explains its deceptively favorable score.
In the second stage, we conduct video-dependent question-answering using SoTA MLLMs on commonsense-related prompts, where models are required to answer based on video observations. Results show that all models—GPT-4o [19], Gemini-2.5-Pro [7], and Qwen-2.5-VL [8]—answered incorrectly (Figure 5), stating the rock reached the ground first, despite the clear visual evidence to the contrary. Moreover, when prompted with commonsense-only questions (i.e., without video context), all models gave the same response, indicating they defaulted to language priors. These results suggest that MLLMs are prone to hallucinations in synthetic video understanding and often rely on priors rather than actual visual input, unlike human perception.
This case study demonstrates that, unlike human perception, current MLLMs struggle to identify abnormalities in synthetic videos that conflict with real-world commonsense, even when the issues are perceptually obvious. The failures stem from limited video perception capabilities and conflicts between language priors and visual observations, often leading to hallucinations. To address this gap, a comprehensive benchmark focused on commonsense and physical reasoning in synthetic video understanding is essential for exploring and guiding MLLMs toward deeper, more reliable visual reasoning.
@@ 3.1 Preliminary
Figure 2: Question Categorization of VideoHallu. We design our benchmark, VideoHallu, with four question categories to probe hallucinations in synthetic video understanding, covering perceptual understanding to abstract reasoning. (a) Alignment checks if the model correctly identifies and understands entities using visual and textual cues. (b) Spatial-temporal Consistency examines whether the model can track entity motion across frames. (c) Common Sense Reasoning tests if the model can reason based on its knowledge. (d) Physics assesses if the model applies physical laws to entity motions and procedural understanding.
3.3 Video Understanding and Evaluation Categorization
We draw inspiration from basic video quality evaluation definitions from MVBench [11] and WorldModelBench [23] to first organize the current challenges of video generations and evaluations in four basic categories (Figure 2). Given the probing target of each question-answering pair and the demand for reasoning abilities or prior knowledge of the LLM backbone to solve the question provided, we divide all question-answering pairs into four major categories with many sub-categories.
Alignment: This question checks whether the model accurately identifies basic entity details and ensures the video content fully aligns with the prompt without omissions or discrepancies.
Spatial-Temporal Consistency: This question evaluates whether the model can detect smooth, consistent changes in objects, actions, and viewpoints over time, without abrupt or abnormal transitions in space or time.
Common Sense Reasoning: This question assesses the model’s ability to apply general knowledge and reasoning to detect conflicts between common sense and the visual context, ensuring it interprets the prompt correctly without misunderstanding or hallucinating entities or actions.
  • Reasoning (CS-R): Tests the model’s ability to interpret problem cues—including emotional or environmental hints—and solve them through reflection and chain-of-thought.
Physics: This question assesses the model’s ability to detect physical inconsistencies, such as violations of gravity, motion dynamics, or conservation laws, requiring careful reasoning about object properties and movements even if not explicitly stated.
  • Motion (P-M): Evaluates the model’s grasp of motion-related physics—like gravity, linear/circular motion, relative movement, and fluid dynamics—spotting inconsistencies or abrupt changes.
Our goal is to move beyond basic metrics like frame consistency or resolution and instead provide a deeper, more rigorous evaluation by identifying visual abnormalities across predefined categories. To this end, we craft targeted adversarial questions designed to explicitly reveal these anomalies. Our core motivation is to assess whether current SoTA MLLMs can effectively detect and interpret such abnormalities—an essential step toward scalable, interpretable video evaluation. We extend these video generation evaluation principles to form our video understanding criteria.
3.4 Data Collection
We first manually crafted 141 adversarial prompts that are challenging for video generation models to generate, including our four predefined categories. For each question-answering pair, each annotator assigns them a super category and a sub category. We gather five expert-level human annotators to craft 3233 question-answering pairs over generated videos, many of which are intentionally designed as reasoning-heavy questions to probe the counter-intuitive or misaligned with the input prompts to test whether video understanding can detect the abnormalities in these videos (Figure 4). [^1] We then do a train/test split on our data with 800 train QA pairs and 1,525 test pairs. We ensure that our train/test split do not have video overlaps. Our questions are meant to challenge current SoTA video generation models that requires understanding of the real world physics and commonsense reasoning abilities, and easily cause unusual phenomenons in the generated videos. For example, the process of a glass breaking, or the process of a bullet shooting into a watermelon are abstract concepts that are challenging for video generation and requiring complex understanding of the real world physics and commonsense reasoning abilities, and easily cause unusual phenomenons in the generated videos. However, being able to detect these abnormalities is a key component for facilitating more robust and interpretable video evaluation, but current SoTA MLLMs struggle at identifying them, which we study in detail in Section 4.

### v2

# VideoHallu: Evaluating and Mitigating Multi-modal Hallucinations on Synthetic Video Understanding

@@ Abstract
Synthetic video generation has gained significant attention for its realism and broad applications, but remains prone to violations of common sense and physical laws. This highlights the need for reliable abnormality detectors that understand such principles and are robust to hallucinations. To address this, we introduce VideoHallu, a benchmark of over 3,000 video QA pairs built from synthetic videos generated by models like Veo2 [1], Sora [2], and Kling [3], paired with expert-crafted counterintuitive QA to evaluate the critical thinking abilities of Multi-modal Large Language Models (MLLMs) on abnormalities that are perceptually obvious to humans but often hallucinated due to language priors. VideoHallu evaluates MLLMs’ abnormality detection abilities with examples across alignment, consistency, commonsense, and physics. We benchmark SOTA MLLMs, including GPT-4o [4], Gemini-2.5-Pro [5], Qwen-2.5-VL [6], Video-R1 [7], and VideoChat-R1 [8]. We observe that these models perform well on many real-world benchmarks like MVBench [9] and MovieChat [10], but still struggle with basic physics-based and commonsense reasoning in synthetic videos. We further show that post-training with Group Relative Policy Optimization (GRPO) [11], using curriculum learning on datasets combining video QA with counterintuitive commonsense and physics reasoning over real and synthetic videos, improves MLLMs’ abnormality detection and critical thinking, demonstrating the value of targeted training for improving their understanding of commonsense and physical laws.
@@ 1 Introduction
Figure 1: Challenges in Abnormality Detection in Synthetic Videos. Synthetic videos often contain counterfactual or commonsense-violating contexts misaligned with reality. Prior score-based evaluations focus on video-prompt alignment and visual consistency, overlooking such abnormalities. VideoHallu addresses this gap by collecting synthetic videos with perceptually obvious abnormalities, paired with crafted questions that probe counterintuitive phenomena or test MLLMs’ critical thinking in detecting such abnormalities. SOTA MLLMs evaluated on VideoHallu show severe hallucinations, indicating these models over-rely on language priors and commonsense knowledge rather than ground their answers in the video content.
Video generation has emerged as a pivotal area of research, attracting significant attention due to its potential to revolutionize AI-Generated Content (AIGC) [12, 13], entertainment [14], and robotics [15]. Despite their visual realism, current video generation models [1, 16, 17] often fail to produce visually grounded contexts aligned with real-world commonsense and physical laws, especially when prompted with out-of-distribution descriptions. This stems from the fact that these models are primarily trained to mimic patterns in the data used for training, without an inherent understanding of the physics-based and commonsense principles that govern the real world. As a result, they can generate plausible-looking, but not physically accurate or semantically consistent videos. Meanwhile, current score-based video evaluators, such as VideoScore [18] and DEVIL [19], primarily focus on frame consistency and visual quality. They tend to ignore perceptually obvious violations of the physical laws in videos generated by Runway Gen2 [17] and Sora [16] (Figure 1), providing limited insight into how well synthetic videos align with reality.
This challenge arises from the misalignment between synthetic videos and reality, underscoring the urgent need for explainable evaluation models [20, 21, 22]. Such evaluation approaches are essential not only for systematically identifying unrealistic generations but also for providing interpretable feedback that can guide the improvement of generative models towards producing reality-consistent content. An emerging approach is to leverage Multi-modal Large Language Models (MLLMs) as evaluators of video content, leveraging their commonsense and physics knowledge by breaking down captions or prompts into questions to assess video quality and alignment [23, 24]. Although QA-based evaluation has been effective for real-world videos, it targets visually grounded content that aligns well with the training data of MLLMs, making the questions relatively easy to answer. In contrast, few studies have explored MLLMs’ performance on synthetic videos, a significantly more challenging domain. As shown in Figure 1, SOTA MLLMs like Gemini-2.5-Pro [5], GPT-4o [25], and Qwen2.5-VL [6] often hallucinate when answering counterintuitive questions about synthetic videos. These models rely on prior knowledge, e.g., “a watermelon should break when shot” in the Counterintuitive Question Answering section of Figure 1, rather than the actual video content, where the watermelon implausibly returns intact after exploding in the Sora-generated video. Moreover, MLLMs struggle with critical thinking in synthetic video contexts. They often fail to detect abnormal physical events and instead generate trivial explanations, such as “the bullet missed” or “it was a toy gun,” as seen in Qwen2.5-VL’s response in the Critical Thinking section of Figure 1.
@@
Motivated by such observations, we propose VideoHallu, a benchmark of expert-crafted, diverse QA pairs covering alignment, consistency, commonsense, and physical reasoning to assess MLLMs’ understanding of synthetically generated videos comprehensively. We evaluate several SOTA models and analyze their limitations in detecting synthetic video abnormalities. Furthermore, to enhance their evaluation capabilities on synthetic videos, we apply curriculum learning [26] and Group Relative Policy Optimization (GRPO) [11] to post-train an MLLM using both real-world (VideoLLaVA [27], PhysBench [28]) and synthetic (VideoHallu) datasets. Our goal is to improve MLLMs’ counterintuitive question answering and critical thinking in commonsense and physical reasoning. Our contributions include:
@@
  • We introduce VideoHallu, a novel benchmark for synthetic video understanding and evaluation, comprising over 3,000 expert-annotated QA pairs spanning alignment, spatial-temporal consistency, commonsense, and physical reasoning across 13 sub-categories.
  • We evaluate state-of-the-art video understanding models on VideoHallu. Top-performing MLLMs (e.g., Qwen-2.5-VL (7B/32B), Gemini-2.5-Pro, GPT-4o) achieve only around 50% accuracy, with severe hallucinations in counterintuitive QA and critical thinking tasks.
@@
  • We fine-tune Qwen2.5-VL-7B using GRPO within a curriculum learning pipeline, leveraging a mix of real-world and synthetic video QA sampled from VideoHallu and other video understanding datasets. Our results show that even a small amount of synthetic data (800 samples) leads to a 3% overall accuracy improvement.
@@ 2 VideoHallu: Benchmark for Synthetic Video Understanding and Evaluation
Preliminary.
We focus on video understanding tasks over synthetic videos, i.e., those generated by video generation models based on text prompts. Ideally, these videos should not only align with the prompts but also reflect real-world plausibility, including: (a) smooth temporal changes in entities, (b) object geometry and motion consistent with commonsense knowledge, and (c) adherence to laws of physics, unless otherwise specified. If a video doesn’t satisfy these criteria, such a video is viewed as an abnormality. To evaluate abnormalities in synthetic videos, we define two types of questions: counter-intuitive and critical thinking. Counter-intuitive questions target abnormal events unlikely in real life (e.g., an exploded watermelon reassembling itself), while critical thinking questions assess whether a model can independently identify visual inconsistencies (e.g., unnatural object breakage). Our goal is to assess whether MLLMs can reason about and evaluate synthetic videos that fall outside the distribution of their training data.
Category Subcategory Description
Alignment Checks model’s accuracy in identifying entities. Entity Counting (A-EC) Quantifies entities present in the scene.
Entity Properties (A-EP) Visual features defining appearance (color, shape, texture).
Entity Recognition & Classification (A-ERAC) Identifies and categorizes entities by attributes.
Spatial Relationships (A-SR) Examines the relative positions of static entities.
Spatial-Temporal Consistency Evaluates abnormal camera/object transitions. Camera Dynamics (SC-CD) Variations in movement, angle, and viewpoint.
Spatial Dynamics (SC-SD) Entity motion, interactions, abrupt changes.
Temporal Dynamics (SC-TD) Tracks appearance shifts and disappearances over time.
Common Sense Reasoning Detects conflicts with visual context. Knowledge (CS-K) Applies general knowledge (geometry, layout, state changes).
Reasoning (CS-R) Interprets cues, enables chain-of-thought solutions.
Physics Detects inconsistencies in physical principles. Conservation (P-C) Mass/energy conservation under external forces.
Constraints & Properties (P-CAP) Understands constraints (rigid bodies, reflections).
Motion (P-M) Physics of movement (gravity, linear/circular motion).
State Transition (P-ST) Physics-driven state changes (heat, phase transitions).
Table 1: Categorization Details of VideoHallu. We design VideoHallu with video QA covering four major categories and multiple sub-categories, enabling rigorous, fine-grained evaluation of visual abnormalities caused by commonsense failures and physical law violations.
Problem formulation.
The motivation of our work comes from the assumption that the language priors within the LLM backbone of the MLLMs may interfere with their understanding of synthetic videos. Our goal is to craft a benchmark with examples composed of synthetic videos with commonsense/physical-law violations that are perceptually obvious to humans but hallucinated by video understanding MLLMs. Let $f_{\mathtt{LVLM}}$, $f_{\mathtt{LLM}}$ denote the LVLM and its LLM backbone, respectively, and $f_{\mathtt{Human}}$ denote the human expert providing ground truth understanding. $f_{\mathtt{LVLM}}(\textit{video},\textit{query})$ can take a video-query pair as input, $f_{\mathtt{LLM}}(\textit{context},\textit{query})$ can take a text-only context-query pair as input, and $f_{\mathtt{Human}}(\textit{context},\textit{query})$ can take multi-modal inputs paired with queries. We denote $\mathcal{V}$ as the set of all contexts within the synthetic video $V$. The context $\mathcal{C}$ denotes the context being probed during the video understanding process, where $\mathcal{Q}$ denotes the query probing this context $\mathcal{C}$. We define a mapping function $\mathit{T}(\cdot)$ that transforms a set of contextual elements into a natural language-formulated text for both the query $\mathcal{Q}$ and context $\mathcal{C}$. This mapping can be performed by either humans or LLMs.
We introduce the contextual distance $d[\cdot,\cdot]$ between two contexts or texts [29]. When two pieces of context convey similar or mutually affirming information, $d$ is small; otherwise, it is large. This distance reflects the alignment between contexts and can be assessed using LLM-as-a-Judge [30, 31, 32] or other model-based evaluators. In the post-training human preference alignment setting, we take the $f_{\mathtt{Human}}(\cdot,\cdot)$ as the ground truth, and ideally, we expect the outputs of both $f_{\mathtt{LVLM}}$, $f_{\mathtt{LLM}}$ are aligned with the human perception and understanding of the real world. The objective function is formulated as follows:
$\max\limits_{V,\mathcal{Q},\mathcal{C}}$ $~{}~{}~{}~{}d[f_{\mathtt{MLLM}}(V,\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(V,\mathit{T}({\mathcal{Q}}))]$ (1)
$s.t.$ $~{}~{}~{}~{}d[f_{\mathtt{LLM}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}}))]\leq\epsilon,$
$~{}~{}~{}~{}d[f_{\mathtt{Human}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(V,\mathit{T}({\mathcal{Q}}))]\geq\delta,~{}\mathcal{C}\subseteq\mathcal{V}$ (2)
The objective function (1) aims to maximize the contextual distance between the answer generated by $f_{\mathtt{MLLM}}$ and the human-annotated ground truth, given a synthetic video $V$ and query $\mathcal{Q}$. The constraints in (2) restrict the problem to a domain where the language prior $f_{\mathtt{LLM}}(\mathit{T}(\mathcal{C}),\mathit{T}(\mathcal{Q}))$, based on the same query $\mathcal{Q}$ and context $\mathcal{C}$, aligns with the ground truth, i.e., the contextual distance is within a tolerance $\epsilon$. At the same time, the video $V$ must contain factual inconsistencies with respect to $\mathcal{C}$ that are detectable by humans when queried by $\mathcal{Q}$, resulting in a contextual distance exceeding a threshold $\delta$. The core context $\mathcal{C}$ must be embedded within the video $V$.
@@
Figure 2: Question Categorization of VideoHallu. We design our benchmark, VideoHallu, with four question categories to probe hallucinations in synthetic video understanding, covering perceptual understanding to abstract reasoning: (a) Alignment checks if the model correctly identifies and understands entities using visual and textual cues. (b) Spatial-temporal Consistency examines whether the model can track entity motion across frames. (c) Common Sense Reasoning tests if the model can reason based on its knowledge. (d) Physics assesses if the model applies physical laws to entity motions and procedural understanding.
Data Collection.
Our data collection pipeline consists of two main stages: The first stage involves generating synthetic videos $V$ with common sense or physics abnormalities, i.e., videos that satisfy constraints (2), where the LLM backbone possesses human-aligned knowledge but MLLMs fail to detect abnormalities, resulting in answers misaligned with human perception. We recruited five human experts to review the defined abnormality categories (detailed in Table 2 and Appendix A) and craft prompts aimed at inducing such abnormalities in generated synthetic videos. In total, they created 141 adversarial prompts, used to generate 987 videos across seven models: Sora [2], Veo2 [1], Runway Gen 2 [17], Kling [3], Pixverse [33], Lavie [34], and CogVideo [35].
In the second stage, we craft adversarial video QA pairs to evaluate MLLMs’ understanding of synthetic videos. Human experts manually review each video $V$ to identify counterintuitive contexts $\mathcal{C}$ that lead to significant discrepancies between MLLM outputs and human perception, i.e., video QA pairs maximizing the objective function (1). They then construct queries $\mathcal{Q}$ as natural language questions $T(\mathcal{Q})$, along with the ground truth answer $f_{\mathtt{Human}}(V,T(\mathcal{Q}))$ based on the identified context. These QA pairs are categorized into sub-categories with definitions provided in Table 2. Each annotator is assigned to write QA pairs highlighting visually clear but semantically abnormal content, difficult for MLLMs to detect. These questions are not designed to trick models, but to probe their ability to catch subtle violations of common sense, physics, or prompt-video mismatches, critical for robust, interpretable video evaluation (Figure 4).  [^1] The dataset contains 3,233 QA pairs, split into 800 for training and 2,433 for testing, with no video overlap between splits.